# Physical modelling of global macrosystems evolution

### v1

@@ Appendix A Data sources
Modis-Kurzweil time series Modis (2003); Kurzweil (2005) considers key evolutionary and technological milestones, focusing on the interplay between demographic growth and technological innovation. They are also referred to as Canonical Milestones. Starting from the Big Bang, it outlines 27 major critical events in the history of the universe, from the formation of galaxies to the rise of the Internet. The time in parentheses refers to the number of years before 2020.
@@
Panov time series Panov (2020); Korotayev and LePoire (2020) represents a historical trajectory of complexity growth, marking key transitions in the development of life and civilization on Earth. They are also referred to as Biological Phase Transitions. This series of 20 events covers critical points in universal and human history, from the origin of life to modern history. Each event signifies a substantial increase in complexity, accelerating over time as each phase becomes shorter. The time in parentheses refers to the number of years before 2020.
@@
The data on GDP growth was obtained from the Maddison Project Database 2020 Bolt and Van Zanden (2020). According to the website, ”The Maddison Project Database provides information on comparative economic growth and income levels over the very long run. The 2020 version of this database covers 169 countries and the period up to 2018”. We aggregated the time series of all $169$ countries to create a global GDP time series, considering data from $700$ CE to $2018$ CE.
@@
The data on U.S. patents was retrieved from the United States Patent and Trademark Office (USPTO) (http://www.uspto.gov/patents/) Lyons (2003). This source provides time series data on all U.S. patents from $1800$ CE to $2020$ CE. Patents are classified by technological codes according to the U.S. Patent Classification System, where each code is represented by an alphanumeric string. From this data, we constructed a time series for both the number of patents and the associated technological codes over time.
@@ Appendix B General Solution of Hyperbolic Differential Equation
In this appendix, we explicitly solve the differential equation governing the models presented in Section III.
@@
$x_{t}=\dfrac{1}{a}\log\left(\dfrac{1}{a(t-t_{0})+e^{-aM_{0}}}\right).$
This expression presents a mathematical singularity at the point $t=t_{0}+a^{-1}e^{-ax_{0}}$.
@@ Appendix C Results of numerical analysis
In this section, we present the results of fitting the models to real data and discuss the meaning of the parameters used. For all analyses, the time variable $t$ is defined as the time before 2020, with the year 2020 corresponding to $t=0$.
For both Biological Phase Transitions and Canonical Milestones, the macrodevelopment rate $\rho_{t}=\frac{dM_{t}}{dt}$ follows the expression (Eq. 5):
$\rho_{t}=\dfrac{1}{e^{-\beta M_{0}}-\beta(t-t_{0})}.$
We fitted $1/\rho_{t}$ against $t$ using a linear model, setting $t_{0}=5$ for Canonical Milestones and $t_{0}=30$ for Biological Phase Transitions. The initial values of $M_{0}$ were taken as $27$ and $20$, respectively. We fitted the parameter $\beta$ and then calculated the macrodevelopment rate $\alpha$ using the relation $\beta=\ln(1+\alpha)$:
$\begin{cases}\beta_{CM}=(0.773\pm 0.009)\times 10^{-10}\\ \alpha_{CM}=(1.16\pm 0.02)\times 10^{-10}\end{cases}$
$\begin{cases}\beta_{BT}=(0.949\pm 0.010)\times 10^{-10}\\ \alpha_{BT}=(1.58\pm 0.03)\times 10^{-10}\end{cases}$
The singularity time $t^{*}$ can be calculated as:
$t^{*}=t_{0}-\frac{1}{\beta}e^{-\beta M_{0}},$
@@
In the case of Biological Phase Transitions, the singularity appears in the past because the time series ends $30$ years before 2020. This highlights the ambiguity and arbitrariness in estimating the exact singularity time, which strongly depends on the chosen endpoint of the dataset. Our results also differ from those reported in Korotayev and LePoire (2020), probably due to different fitting techniques. Additionally, the quite large error margins prevent a precise estimate of the singularity year.
@@
The macrodevelopment parameter $\alpha$ represents the combinatorial success rate, quantifying the fraction of potential combinations that actually produce successful innovations. The two parameters $\alpha_{CM}$ and $\alpha_{BT}$ are close in magnitude, reflecting the same underlying process, irrespective of the specific events considered in the two time series. From the parameters, we can estimate the number of expected new Canonical Milestones in the future. From the discrete version of the model:
$M_{t+1}-M_{t}=\sum_{i=1}^{M_{t}}\alpha^{i}\binom{M_{t}}{i},$
we can compute the discrete macrodevelopment rate $\rho_{t}=M_{t+1}-M_{t}$. For $M_{t}=27$, this is approximately $\rho_{t}\approx 0.10$. This implies that the next Canonical Milestone is expected to occur about $101/\rho_{t}=$ 10 years after $t_{0}=2015$, around the year 2025.
@@
After this point, with $M_{t}=28$, the rate increases to $\rho_{t}\approx 0.20$, meaning the next milestone is expected in the next $5$ years, thus around 2030. Notably, this prediction bypasses the singularity at 2027, emphasizing that in the discrete version of the model, the concept of a singularity does not apply. The discrete formulation thus allows for more accurate forecasting of future events, overcoming the limitations of continuous models.
@@
$N_{t}=\dfrac{1}{N_{0}^{-1}-\gamma(t-t_{0})},$
and fitting $1/N_{t}$ against time, imposing $t_{0}=0$ and $N_{0}\approx 7.79\cdot 10^{8}$, we obtain:
$\gamma=(1.61\pm 0.06)\times 10^{-12},$
$t^{*}_{WP}=\dfrac{1}{N_{0}\gamma}=-34.6\pm 0.6\ (2054.6).$
Note the relatively small error in the singularity estimate, which is due to the shorter time interval considered for this dataset compared to the previous ones, reducing uncertainty. Once again, our predictions differ from previous literature due to the use of different data and time periods.
The macrodevelopment parameter $\gamma$ can be interpreted as a birth rate, indicating the fraction of potential interactions between individuals that lead to successful reproduction. Considering the discrete version of the model:
@@
we find, for example, that with the current population size of approximately $8\cdot 10^{9}$, the predicted increase in the next year is about $\gamma N_{t}^{2}\approx 1\cdot 10^{8}$. This estimate aligns well with the actual population growth observed between 2020 and 2023, which averaged around $0.75\cdot 10^{8}$ per year.
@@
$K_{t}\approx\dfrac{1}{\left(K_{0}^{-1/2}-\delta(t-t_{0})\right)^{2}}.$
By fitting the linear relation between $1/K_{t}^{2}$ and time, and using the initial conditions $t_{0}=0$ and $K_{0}=3.402405\cdot 10^{6}$, we obtain:
$\delta=(1.96\pm 0.03)\times 10^{-5},$
$t^{*}_{GDP}=\dfrac{1}{\sqrt{K_{0}}\delta}=-26.7\pm 0.4\ (2046.7).$
The macrodevelopment parameter $\delta$ represents the rate of GDP growth. Using the discrete dynamics:
@@
we estimate that in the year following 2020, GDP should grow by approximately $\delta K_{t}^{3/2}\approx 1.2\cdot 10^{5}$, consistent with the growth rate measured in the last three years (2020-2023), which was about $3\%$ per year.
@@
From the exponential fit, and setting $C_{0}$ and $P_{0}$, we find $\beta=(4.18\pm 0.02)\times 10^{-5}$. In this case, there is no singularity, even in the continuous version of the model.

### v2

@@ Appendix A Data sources
Modis-Kurzweil time series modis2003limits ; kurzweil2005singularity considers key evolutionary and technological milestones, focusing on the interplay between demographic growth and technological innovation. They are also referred to as Canonical Milestones. Starting from the Big Bang, it outlines 27 major critical events in the history of the universe, from the formation of galaxies to the rise of the Internet. The time in parentheses refers to the number of years before 2020.
@@
Panov time series panov2020singularity ; korotayev202021st represents a historical trajectory of complexity growth, marking key transitions in the development of life and civilization on Earth. They are also referred to as Biological Phase Transitions. This series of 20 events covers critical points in universal and human history, from the origin of life to modern history. Each event signifies a substantial increase in complexity, accelerating over time as each phase becomes shorter. The time in parentheses refers to the number of years before 2020.
@@
The data on GDP growth was obtained from the Maddison Project Database 2020 bolt2020maddison . According to the website, ”The Maddison Project Database provides information on comparative economic growth and income levels over the very long run. The 2020 version of this database covers 169 countries and the period up to 2018”. We aggregated the time series of all $169$ countries to create a global GDP time series, considering data from $700$ CE to $2018$ CE.
@@
The data on U.S. patents was retrieved from the United States Patent and Trademark Office (USPTO) (http://www.uspto.gov/patents/) lyons2003united . This source provides time series data on all U.S. patents from $1800$ CE to $2020$ CE. Patents are classified by technological codes according to the U.S. Patent Classification System, where each code is represented by an alphanumeric string. From this data, we constructed a time series for both the number of patents and the associated technological codes over time.
@@ Appendix B General Solution of Hyperbolic Differential Equation
In this appendix, we explicitly solve the differential equation governing the models presented in Section III. Throughout the following derivation, time is expressed in years $[y]$.
@@
$x_{t}=\dfrac{1}{b}\ln\left(\dfrac{1}{e^{-bx_{0}}-ab(t-t_{0})}\right).$
This expression presents a mathematical singularity at the point $t=t_{0}+a^{-1}b^{-1}e^{-bx_{0}}$. Note that Eq. (5) is derived from:
$\dfrac{dM_{t}}{dt}\approx\overline{\nu}\alpha 2^{M_{t}}=\overline{\nu}\alpha e^{\ln 2M_{t}}$
with $a=\overline{\nu}\alpha$ and $b=\ln 2$.
@@ Appendix C Results of numerical analysis
In this section, we present the results of fitting the models to real data and discuss the meaning of the parameters used. For all analyses, the time variable $t$ is defined as the time before 2020, i.e., $\tau=2020-t$, with the year 2020 corresponding to $\tau=0$.
For both Biological Phase Transitions and Canonical Milestones, the macrodevelopment rate $\rho_{t}=\frac{dM_{t}}{dt}$ follows the expression (Eq. (8)):
$\rho_{t}=\dfrac{1}{\rho_{0}^{-1}-\beta(t-t_{0})}.$
With the time-shift $\tau=2020-t$, as also used in Figure 1, the macrodevelopment rate reads:
$\rho_{\tau}=\dfrac{1}{\rho_{0}^{-1}-\beta(2020-\tau-t_{0})}=\dfrac{1}{\rho_{0}^{-1}+\beta(\tau-\tau_{0})},$
where $\tau_{0}=2020-t_{0}$ is the shifted initial condition. This expression was used in the fit.
We fitted $1/\rho_{\tau}$ against $\tau$ using a linear model, setting $\tau_{0}=50$ ($t_{0}=1970$) for Canonical Milestones and $\tau_{0}=70$ ($t_{0}=1950$) for Biological Phase Transitions. The initial values of $\rho_{0}$ were taken as $\rho_{0}^{CM}=0.0222$ and $\rho_{0}^{BT}=0.0250$ from the time series, respectively.
We fitted the parameter $\beta$ and then calculated the macrodevelopment rate $\alpha$ using the relation $\beta=\ln(1+\alpha)$:
$\begin{cases}\beta_{CM}=0.513\pm 0.007\\ \alpha_{CM}=0.670\pm 0.012\end{cases}$
$\begin{cases}\beta_{BT}=0.608\pm 0.002\\ \alpha_{BT}=0.837\pm 0.004\end{cases}$
The singularity time $\tau^{*}$ can be calculated as:
$\tau^{*}=\tau_{0}-\frac{1}{\rho_{0}\beta}$
@@
Note that, from our definition of $\tau$, its negative values correspond to times after 2020 CE, while positive values indicate times before 2020 CE. In the case of Biological Phase Transitions, the singularity appears in the past because the time series ends $70$ years before 2020. This highlights the ambiguity and arbitrariness in estimating the exact singularity time, which strongly depends on the chosen endpoint of the dataset. Our results also differ from those reported in korotayev202021st , probably due to different fitting techniques. Additionally, the quite large error margins prevent a precise estimate of the singularity year.
@@
The macrodevelopment parameter $\alpha$ represents the combinatorial success rate, quantifying the fraction of potential combinations that actually produce successful innovations. The two parameters $\alpha_{CM}$ and $\alpha_{BT}$ are close in magnitude, suggesting a similar underlying process, irrespective of the specific events considered in the two time series.
Regarding the number of events $M_{t}$, the continuous dynamics described in equation (9), with parameter $\beta$ obtained by fitting the macrodevelopment rate, may introduce errors when the time intervals between events are large. This is particularly relevant in the case of Canonical Milestones and Biological Phase Transitions, where we have only $27$ or $20$ data points over a time range of $10^{10}$ years. In Eq. (9), $M_{t}$ is computed by a continuous integration of $\rho_{t}$, which can lead to discrepancies compared to the actual discrete dynamics when the time intervals are large. This issue does not arise for world population, GDP, and patents, where denser data sets minimize errors in the transition from discrete to continuous models.
To avoid these errors, the sequence of events $M_{t}$ can be reconstructed directly using the discrete model based on $\rho_{t}$:
$\begin{cases}\rho_{t+\Delta t}=\rho_{t}+\beta\rho_{t}^{2}\Delta t,\\ M_{t+\Delta t}=M_{t}+\rho_{t}\Delta t.\end{cases}$ (17)
Here, the first equation discretizes equation (7), while the second reproduces the sequence $M_{t}$ shown in Figure 1. This ensures that the dynamics remain faithful to the discrete model, particularly when the time intervals between events are large, as is the case for milestones.
Using the discrete model, we can estimate the timing of new Canonical Milestones in the short-term future. It is important to note that the continuous version of the model provides an exact analytical continuation of the discrete one only far from the singularity. Parameters obtained by fitting the continuous model can reliably predict $\rho_{t}$ over short time intervals following the last observed event. To extend the prediction range, the sequence might be updated with new data points (either real or estimated), and the parameters should be adjusted based on the updated sequence.
From the second equation in (17), we can compute the macrodevelopment rate $\rho_{t}$ at the next stage of the evolution and subsequently estimate the number of milestones $M_{t}$. Let’s focus on Canonical Milestones. The inverse of $\rho_{t}$ represents the inter-time, i.e., the time interval between two consecutive events. Starting with $\rho_{0}=0.022$ at $\tau_{0}=50$, the next event $M_{t}=27$ is expected to occur after $1/\rho_{\tau}=\Delta\tau\approx 45$ years. This aligns with the last point of the time series, corresponding to $\tau=5$ years before 2020. At that time, the rate will be:
$\rho_{\tau}=\rho_{0}+\beta\rho_{0}^{2}\Delta\tau\approx 0.0336\ [y^{-1}].$
@@
Thus, the next Canonical Milestone $M_{t}=28$ is expected approximately $1/\rho_{\tau}\approx 30$ years after $\tau=5$, corresponding to $t=2045$ CE. Iterating the procedure, the following event $M_{t}=29$ is predicted to occur around $t=2065$ CE. This last prediction is purely speculative. As discussed earlier, predictions far from the last observed point would require updating the parameters using new data. Nevertheless, the prediction for 2065 bypasses the singularity at 2057, illustrating that in the discrete model, the concept of a singularity does not apply. The discrete formulation thus enables forecasting of future events beyond the singularity, addressing the limitations of continuous models.
As a final remark, note that the scaling factor $\overline{\nu}$ in Eq. (6) was not explicitly considered, as we worked directly with the macrodevelopment rate, which does not depend on this factor. In fact, any rescaling of time affects the rate in the same way, ensuring that the results remain unchanged by the choice of $\overline{\nu}$. Nevertheless, its value can still be determined from the data using the relation $\rho_{t}=\overline{\nu}e^{\beta M_{t}}$, which directly leads to $\overline{\nu}=\rho_{t}e^{-\beta M_{t}}$. By considering the first point in the time series to minimize discretisation errors, we obtain $\overline{\nu}\approx 10^{-10}\ [y^{-1}]$ for CM and $\overline{\nu}\approx 2.2\cdot 10^{-10}\ [y{-1}]$ for BT. When modelling the number of milestones, the time-scaling factor $\overline{\nu}$ becomes necessary, since the right-hand side of Eq. (6) is always greater than 1, while the actual time scale requires a rate smaller than 1.
@@
$N_{t}=\dfrac{1}{N_{0}^{-1}-\nu_{\gamma}(t-t_{0})},$
we shift the time variable to obtain:
$N_{\tau}=\dfrac{1}{N_{0}^{-1}+\nu_{\gamma}(\tau-\tau_{0})},$
where $\tau=2020-t$ and $\tau_{0}=2020-t_{0}$.
We fit $1/N_{\tau}$ against time, imposing $\tau_{0}=0$ ($t_{0}=2020$) and $N_{0}\approx 7.79\cdot 10^{9}$, we obtain:
$\nu_{\gamma}=(2.01\pm 0.06)\times 10^{-12}\ [y^{-1}],$
$\tau^{*}_{WP}=-\dfrac{1}{N_{0}\nu_{\gamma}}=-63.8\pm 0.6\ [y],\ (t^{*}_{WP}\approx 2084\ [y]).$
Note the relatively small error in the singularity estimate, which is due to the shorter time interval considered for this dataset compared to the previous ones, as well as the larger number of points, reducing uncertainty. Once again, our predictions differ from previous literature due to the use of different data and time periods.
Due to the smaller time scale considered, in this case we can safely put $\overline{\nu}=1\ [y^{-1}]$ cortes2022biocosmology . As a consequence, the macrodevelopment parameter $\gamma$ is equal to $\nu_{\gamma}$, without dimension. This parameter can be interpreted as a birth rate, indicating the fraction of potential interactions between individuals that lead to successful reproduction. Considering the discrete version of the model:
@@
we find, for example, that with the current population size of approximately $8\cdot 10^{9}$, the predicted increase in the next year is about $\gamma N_{t}^{2}\approx 1.2\cdot 10^{8}$. This estimate aligns with the actual population growth observed between 2020 and 2023, which averaged around $0.75\cdot 10^{8}$ per year.
@@
$K_{\tau}\approx\dfrac{1}{\left(K_{0}^{-1/2}+\nu_{\delta}(\tau-\tau_{0})\right)^{2}}.$
By fitting the linear relation between $1/K_{\tau}^{2}$ and time, and using the initial conditions $\tau_{0}=0$ ($t_{0}=2018$) and $K_{0}=3.402405\cdot 10^{6}$, we obtain:
$\nu_{\delta}=(1.96\pm 0.03)\times 10^{-5}\ [y^{-1}],$
$\tau^{*}_{GDP}=-\dfrac{1}{\sqrt{K_{0}}\nu_{\delta}}=-27.7\pm 0.4\ [y],\ (t_{GDP}^{*}\approx 2048\ [y]).$
As for the world population, in this case the time scaling factor can be put equal to $\overline{\nu}=1\ [y^{-1}]$. The dimensionless macrodevelopment parameter $\delta=(1.96\pm 0.03)\times 10^{-5}$ represents the rate of GDP growth. Using the discrete dynamics:
@@
we estimate that in the year following 2018, GDP should grow by approximately $\delta K_{t}^{3/2}\approx 1.2\cdot 10^{5}$ billions of USD dollars. The actual GDP growth observed over the years (2019-2022) was about 3% per year. Given that $K_{t}\approx 3.40\times 10^{6}$ billions of USD dollars at $t_{0}=2018$, this corresponds to an annual increase of approximately $1.0\times 10^{5}$ billions of USD dollars, aligning well with our estimate.
@@
The equation for patents (Eq. (15)) is dimensionless, so in this case, there is no need to account for a dimensional factor. From the exponential fit, and setting $C_{0}$ and $P_{0}$, we find $\beta=(4.18\pm 0.02)\times 10^{-5}$. In this case, there is no singularity, even in the continuous version of the model.